OpenStudy (2yoututu11):

A triangle has a base of 9 inches and a height of 4 inches. What is the area of the triangle? A. 13 square inches B. 15 square inches C. 18 square inches D. 36 square inches

OpenStudy (anonymous):

1/2 Bh

OpenStudy (albert0898):

Area of a triangle = base * height * 1/2 OR base * height divided by 2 Area = 9 * 4 divided by 2 Area = 36 / 2 Area = 18 Answer = [C] 18 Square Inches

OpenStudy (anonymous):

it was c the 18 sq inches. 1/2(9*4)